OpenStudy (anonymous):

(1.8 x 10^1)(7 x 10^5)

OpenStudy (anonymous):

it look like you are foiling

OpenStudy (anonymous):

1,260,000

OpenStudy (anonymous):

i know tht but what about ina simplified form

OpenStudy (anonymous):

like for example 4.1 x 10^-7

OpenStudy (e.mccormick):

\((1.8 \times 10^1)(7 \times 10^5) \) Rules for order of operations allows: \(1.8 \times 7 \times 10^1\times 10^5 \) Now multiply the 1.8 and 7, then use exponent rules on the 10s to a power.